OpenStudy (anonymous):

Evaluate the difference quotient and simplify the result. f(x) 1 over x+2.. (1/x+2) then below this function i have f(x+ delta x) - f(x) /over delta x. My original answer of 1/(x+2)(x + delta x+2) is wrong. Can someone please tell me where I am wrong?

OpenStudy (anonymous):

so \[f \left( x \right)=\frac{ 1 }{ x+2 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \[\frac{ \Large\frac{ 1 }{ (x+h)+2 }-\frac{ 1 }{ x+2 } }{ h }\]is the difference quotient, yes?

OpenStudy (anonymous):